# The Surprising Soupability of Documents in State Space Models

### v1

@@ A.4 Decoding Hyperparameters
Finetuned on $n=0\ gold+0\ distractors$ (QA-only)
@@
Finetuned on $n=2\ gold+0\ distractors$ documents
@@
[…]
Finetuned on $n=2\ gold+3\ distractors$ documents
@@
[…]
Finetuned on $n=2\ gold+5\ distractors$ documents
@@
[…]
Finetuned on $n=2\ gold+8\ distractors$ documents
@@
A.4.1 Compute Resources
@@ Appendix B Additional Results and Analysis
B.1 HotPotQA Extended Analysis
B.1.1 Mamba2-2.7B Results
@@ B.1 HotPotQA Extended Analysis
B.1.2 Mamba2-8B Results
Table 7 expands on the main paper’s Table 1 by providing a more detailed view of HotpotQA results for the Mamba2-8B model trained on 5 documents (2 gold + 3 distractors). At inference time, we evaluate the same checkpoint across a range of input sizes, including a no-context setting (query-only) and settings with 2 to 10 documents, where each includes 2 gold documents and $n{-}2$ distractors.
@@ B.1.2 Mamba2-8B Results
Table 7: HotpotQA performance (Exact Match / F1) for Mamba-2 8B trained on 5 documents (2 gold + 3 distractors) and tested on $n$ documents, with $2$ gold and $(n-2)$ distractors. We compare pretrained models, decoder-only fine-tuning, and encoder-decoder fine-tuning, under different pooling operators (Avg, Sum, Max) and normalization settings. Underline marks models tested on the same docs they are trained on.
Method Test on $2$ gold + $(n-2)$ distractors
@@
[…]
@@
[…]

### v2

@@ A.5 Decoding Hyperparameters
Finetuned on n=0gold+0distractorsn=0\ gold+0\ distractors (QA-only)
@@
Finetuned on n=2gold+0distractorsn=2\ gold+0\ distractors documents
@@
Soup w/ Average 47.4 / 61.0 39.9 / 52.7 33.0 / 45.5 28.8 / 40.4 25.5 / 36.8 23.8 / 34.2 21.9 / 31.8 19.7 / 29.2 19.0 / 28.2
Finetuned on n=2gold+3distractorsn=2\ gold+3\ distractors documents
@@
Soup w/ Average 49.1 / 62.9 44.8 / 58.3 41.5 / 54.4 38.4 / 51.0 35.7 / 48.2 33.7 / 45.5 32.7 / 43.8 30.1 / 41.3 28.6 / 39.3
Finetuned on n=2gold+5distractorsn=2\ gold+5\ distractors documents
@@
Soup w/ Average 48.8 / 63.0 44.2 / 57.8 40.3 / 53.5 37.7 / 50.5 36.0 / 48.2 33.6 / 45.6 32.0 / 43.7 30.1 / 41.2 28.5 / 39.5
Finetuned on n=2gold+8distractorsn=2\ gold+8\ distractors documents
@@
A.5.1 Transformer Cache Souping Experimental Setup
To test the viability of extending our souping method to Transformer architectures, we conducted a comparative experiment using LLaMA3-8B Instruct. We finetuned the model for one epoch on 10,000 examples from the HotpotQA dataset, where each example consisted of two gold documents and eight distractor documents. We compared two distinct fine-tuning methodologies:
  • Normal Fine-tuning (Baseline): The model was fine-tuned on the simple concatenation of all 10 documents and the question. We used a learning rate of $1\times 10^{-5}$ for this setup.
  • KV Cache Souping Fine-tuning: To handle variable-length inputs, all documents were left-padded to match the longest document in the batch. Each document was then processed independently to generate its key-value (KV) cache. The resulting caches were averaged to create a single, souped KV cache, which was then used alongside the question for downstream reasoning. A learning rate of $1\times 10^{-6}$ was used for this approach.
A.5.2 Compute Resources
@@ Appendix B Additional Results and Analysis
B.1 Scalability of Soup Sizes
To further probe the limits of scalability, we conducted experiments up to 256 documents. To stay consistent with previous results for fair comparison, we evenly splitted the original 8192 sequence length NIAH tasks into 64, 128, and 256 segments. The full results of this scalability analysis are presented in Table 8. We first finetuned a Mamba2-8B model on the NIAH dataset using 64 segments for 25,000 samples. While this model generalized well to nearby segment counts, its performance dropped sharply when tested on 128 or 256 segments. However, by continuing to finetune this checkpoint on 128 segments for 8,400 samples, the model’s performance at these higher counts improved dramatically. Compared to concatenating finetuining baseline’s result 35.55 in Table 3, the continued checkpoint reaches a score of 39.00 at 128 segments and 37.25 at 256 segments, outperforming the baseline while maintaining robustness across smaller soup sizes. This two-stage process demonstrates that while extrapolation is challenging, the model can be effectively trained to handle very large soup sizes when provided with appropriate data.
Test Segments
Finetuned on 64 Segments
Continued on 128 Segments
2 28.65 23.65
4 34.40 26.60
8 37.95 27.70
16 40.05 30.25
32 41.65 32.30
64 39.75 37.80
128 15.15 39.00
256 0.10 37.25
Table 8: Scalability results (EM %) on the NIAH dataset for a Mamba2-8B model with an 8k sequence length. We compare a model finetuned on 64 segments against a model that was subsequently trained further on 128 segments. Continued training significantly improves performance on larger soup sizes, demonstrating the model’s capacity to adapt to wider contexts.
B.2 HotPotQA Extended Analysis
B.2.1 Mamba2-2.7B Results
@@ B.2 HotPotQA Extended Analysis
B.2.2 Mamba2-8B Results
Table 9 expands on the main paper’s Table 1 by providing a more detailed view of HotpotQA results for the Mamba2-8B model trained on 5 documents (2 gold + 3 distractors). At inference time, we evaluate the same checkpoint across a range of input sizes, including a no-context setting (query-only) and settings with 2 to 10 documents, where each includes 2 gold documents and n2n{-}2 distractors.
@@ B.2.2 Mamba2-8B Results
Table 9: HotpotQA performance (Exact Match / F1) for Mamba-2 8B trained on 5 documents (2 gold + 3 distractors) and tested on $n$ documents, with 22 gold and (n2)(n-2) distractors. We compare pretrained models, decoder-only fine-tuning, and encoder-decoder fine-tuning, under different pooling operators (Avg, Sum, Max) and normalization settings. Underline marks models tested on the same docs they are trained on.
Method Test on $2$ gold + (n2)(n-2) distractors
@@
Average + Norm Before 35.9 / 47.8 50.7 / 63.9 49.7 / 63.0 47.8 / 60.9 45.7 / 58.5 43.3 / 55.9 41.8 / 54.2 39.6 / 51.8 38.1 / 50.2
Average + Norm After 50.4 / 65.2 47.6 / 61.1 45.3 / 58.8 42.3 / 55.6 40.3 / 53.2 38.1 / 50.7 36.2 / 48.8 34.9 / 46.8 33.4 / 44.9
Average + Norm Before & After 6.9 / 10.7 27.7 / 36.8 41.6 / 54.0 42.2 / 54.7 40.5 / 52.8 38.7 / 51.0 36.5 / 48.8 35.2 / 47.5 33.8 / 45.6
@@
Sum + Norm Before 8.8 / 13.6 10.3 / 15.0 9.1 / 13.9 8.1 / 13.0 6.9 / 11.9 6.3 / 11.2 5.8 / 10.6 5.2 / 10.0 4.6 / 9.4
Sum + Norm After 51.9 / 66.1 49.7 / 63.2 46.9 / 60.2 43.6 / 56.6 41.5 / 54.2 39.8 / 52.0 38.4 / 50.7 36.6 / 48.6 34.7 / 46.2
Sum + Norm Before & After 10.2 / 16.3 32.0 / 42.7 40.3 / 53.1 40.3 / 52.9 38.4 / 50.9 37.3 / 49.8 35.8 / 48.2 34.6 / 46.6 33.2 / 44.8
@@ B.4 Ruler QA_1
Appendix C Use of Large Language Models
Large language models (LLMs) were utilized to assist in the preparation of this manuscript in two capacities. First, LLMs served as a writing aid for copy-editing and for improving the clarity and readability of the text. The authors drafted all original content and retained full editorial control, evaluating all suggestions to ensure they aligned with our intended meaning. Second, LLMs were used to support the literature discovery process. The authors provided seed papers on relevant topics and leveraged the models to find additional related work. All papers suggested by the LLMs were subsequently vetted by the authors for relevance before inclusion.